Grade 11General Physics

ABC is an isosceles triangle right angled at a. Forces of agnitude 2∫2,5,6N act along BC,CA,AB respectively.the magnitide of their resltant force in N is

SUJAL CHOUDHARY
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
As the triangle is isosceles and right calculated at A 
Rcosθ=6cos0∘+22–√cos(180∘−B)+5cos270∘Rcos⁡θ=6cos⁡0∘+22cos⁡(180∘−B)+5cos⁡270∘ 
=> 6−22–√cosB6−22cos⁡B 
Rsinθ=6sin0+22–√sin(180∘−B)+5sin270∘Rsin⁡θ=6sin⁡0+22sin⁡(180∘−B)+5sin⁡270∘ 
=22–√sinB−5=22sin⁡B−5 
∴R2=36+8cos2B−242–√cosB+8sin2B+25−202–√sinB∴R2=36+8cos2⁡B−242cos⁡B+8sin2⁡B+25−202sin⁡B 
=44−242–√cos45∘−202–√sin45∘+25=44−242cos⁡45∘−202sin⁡45∘+25 
[ABC is a privilege ∠d∠d isosceles triangle ] 
∴[∠B=∠C=45∘]∴[∠B=∠C=45∘] 
=> 69−242–√12–√=69−24212=202–√12–√20212 
=> 69−24−2069−24−20 
=> 2525 
=> R=5